[Brahms-l] Shift report 20050606 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Jun 06 2005 - 09:31:30 EDT
Shifter leader was sl
Shifters were gl


1.RHIC working well. Long store starting at 2 am.

2.Done sweep of mrs B 4KG angles. Mrs at 90 deg full, FFS at 3 deg, BFS at 2.3 deg. Full field.

3. T3 A3-xu channel bad.
H2s scaler signal at 0  since about 7 o'clock and Trigger 2 gone. HV ok. 
Tried restarting VME crate, but without any effect.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 6 09:31:52 2005

This archive was generated by hypermail 2.1.8 : Mon Jun 06 2005 - 09:32:04 EDT